OpenStudy (anonymous):

Simplify the expression using the properties of exponents: (x^-2y^0)^2(4x^3y^-3)^3 ______________________________ (3x^0y^2)^2(x^-1y^3)^2

OpenStudy (anonymous):

ignore all variable with power of 0

OpenStudy (anonymous):

what imranmeah said

OpenStudy (anonymous):

again:for complex expression please use latex (i.e. the equation editor)

OpenStudy (anonymous):

\[\frac{(x^{-2})^2(4x^3y^-3)^3 }{(3y^2)^2(x^{-1}y^3)^2}\]

OpenStudy (anonymous):

Where is the equation editor?

OpenStudy (anonymous):

\[\frac{x^{-4}4^3x^9y^{-9}}{9y^4x^{-2}y^6}\]

OpenStudy (anonymous):

\[\frac{4^3x^5y^{-9}}{9y^{10}x^{-2}}\]

OpenStudy (anonymous):

\[\frac{64x^7}{9y^{19}}\] although i certainly could have made a mistake along the way

OpenStudy (anonymous):

i didn't see a mistake so I am agreeing with that answer.